# ODTFormer: Efficient Obstacle Detection and Tracking with Stereo Cameras Based on Transformer

### v2

@@ Abstract
Obstacle detection and tracking represent a critical component in robot autonomous navigation. In this paper, we propose ODTFormer, a Transformer-based model to address both obstacle detection and tracking problems. For the detection task, our approach leverages deformable attention to construct a 3D cost volume, which is decoded progressively in the form of voxel occupancy grids. We further track the obstacles by matching the voxels between consecutive frames. The entire model can be optimized in an end-to-end manner. Through extensive experiments on DrivingStereo and KITTI benchmarks, our model achieves state-of-the-art performance in the obstacle detection task. We also report comparable accuracy to state-of-the-art obstacle tracking models while requiring only a fraction of their computation cost, typically ten-fold to twenty-fold less. The code and model weights will be publicly released.
@@ I Introduction
In this paper, we present ODTFormer, a Transformer-based [23] model, to address the aforementioned obstacle detection and tracking challenges. Unlike existing stereo-based models, which construct a pixel-wise cost volume for disparity matching [24, 25, 26, 27, 14], we propose to use deformable cross-attention [28] from 3D voxel queries to 2D stereo image features to compute the matching cost. Compared with the pixel-wise cost volume used in [14], we directly construct it in the 3D space, conforming better to the scene geometry. More importantly, our approach disentangles dataset specifics from the model design and thus shows better generalization than [14]. The cost volume is then processed by a U-Net decoder to produce occupancy voxel grids in a progressive manner for efficiency purposes.
@@
To validate the effectiveness of our approach, we conduct comprehensive evaluations on DrivingStereo [29] and KITTI [30] datasets. In the obstacle detection task, we demonstrate a significant performance improvement compared to prior works [24, 25, 26, 27, 14]. Additionally, our approach exhibits greater generalization across various camera parameters and resolutions than [14]. In the obstacle tracking task, we compare against scene flow-based approaches and show that our approach can achieve comparable results to state-of-the-art methods while needing much lighter computation cost. Detailed ablation studies are conducted to validate different design choices. In summary, our contributions are:
@@ II Related Works
Figure 2: Illustration of the overall architecture design. Left: For obstacle detection, we first extract multi-scale 2D feature maps [31, 32] for each of the stereo images. We then encode the voxels in the ROI to cross-attend to the image features to compute the matching cost, which is refined by integrating the matching cost from a finer granularity. Such a cost volume is directly constructed in the 3D space, which conforms better to the scene geometry, disentangles dataset specifics from model design, and thus generalizes well. It is then progressively decoded into occupancy voxel grids. Right: For obstacle tracking, we cast it as a matching problem by finding the correspondences of voxels across two consecutive frames, where we incorporate physical constraints to improve both the accuracy and efficiency. Both the detection and tracking modules can be jointly optimized in an end-to-end manner and run efficiently.
@@ II-B Obstacle Perception
Obstacle tracking aims at estimating the motion of obstacles across time. While most of the previous works are based on LiDAR point cloud [36, 37] or camera-LiDAR fusion [38], visual methods, especially stereo-based ones, remain under-explored. Existing visual approaches rely on object detectors or Kalman filters [10, 12, 11], limiting generalizability and requiring meticulous parameter tuning. We introduce voxel flow, which is the 3D grid-to-grid voxel-level motion of occupied voxels while achieving near real-time obstacle tracking.
In parallel to our work, there exist seminal occupancy detection approaches tailored for autonomous driving [8, 7]. These approaches commonly operate under the assumption of having access to surrounding view images. In contrast, our approach utilizes only a pair of stereo images and emphasizes computational efficiency.
@@ III-A2 Cost volume construction via cross-attention
A crucial difference of our proposed approach from [14] is our novel design of the cost volume construction module, where we employ cross-attention from 3D voxel queries to 2D stereo image features. Compared with the pixel-wise cost volume used in [14], we directly construct it in the 3D space, conforming better to the scene geometry. More importantly, our approach disentangles dataset specifics from the model design and thus shows better generalization than [14]. We will introduce this module in detail in Sec. III-B.
@@ III-B Cost Volume Construction via Cross-Attention
Specifically, for the voxel centered at $p$, we first sample points on multi-scale feature maps $({\mathbf{F}}_{i}^{L},{\mathbf{F}}_{i}^{R})$ (similar to the pixel-aligned features ${\mathbf{f}}$ but on multi-scale features instead of just the last level) around the projected points of the voxel centroid $p$ based on the offset $\delta$, respectively. $\delta$ is generated by a learnable offset sampler $g$ [28] based on the voxel’s encoding ${\mathbf{q}}$, $\delta=g({\mathbf{q}})$, where $g$ is an MLP plus a sigmoid output layer (scaled by a local window size of $32\times 32$ pixels). We then directly concatenate the sampled features along the channel dimension and employ a scale-specific MLP to compute the cross-view matching cost ${\mathbf{c}}_{i}$:
${\mathbf{c}}_{i}(p,{\mathbf{q}},{\mathbf{F}}_{i})=\texttt{MLP}_{i}($ ${\mathbf{F}}_{i}^{L}(\mathcal{P}(p,\theta^{L})+\delta)\oplus$ (3)
${\mathbf{F}}_{i}^{R}(\mathcal{P}(p,\theta^{R})+\delta)),$
where $\oplus$ denotes the feature concatenation. $\texttt{MLP}_{i}$ has two linear layers with batch normalization and ReLU activation in between. Instead of using a single offset, multiple ones are usually generated to enhance the computing of the matching cost, and different matching costs for different samples are aggregated as follows.
${\mathbf{C}}_{B}(p,{\mathbf{q}},{\mathbf{F}})=\sum_{i\in\{1,2,3,4\}}\sum_{s=1}^{N_{s}}\mathbf{A}_{is}\mathbf{W}_{s}{\mathbf{c}}_{i}(p,\delta_{s},{\mathbf{F}}_{i}),$ (4)
where $N_{s}$ is the total number of sampling points. $\delta_{s}$ denotes the $s$-th learned offset. $\mathbf{A}_{is}\in[0,1]$ is the learnable weights for the cost generation and $\mathbf{W}_{s}\in\mathbb{R}^{D\times D}$ are learnable parameters. We can get a cost volume ${\mathbf{C}}_{B}\in\mathbb{R}^{n_{x}\times n_{y}\times n_{z}\times D}$ by computing the matching cost for all voxels in ${\mathbf{Q}}_{B}$.
Deformable cross-attention has been used in previous works [6, 7, 8, 9] to aggregate multi-view features. But instead of simply averaging them as in existing work, we leverage the geometric constraints between two stereo images and use the deformable attention to construct a cost volume in 3D, leading to better accuracy, as evidenced by our ablation studies.
Matching cost refine. We further refine our cost volume by combining information at various granularities. Specifically, we partition the ROI into ${\mathbf{Q}}_{S}$ with half resolution of ${\mathbf{Q}}_{B}$ in the dimension $\frac{n_{x}}{2}\times\frac{n_{y}}{2}\times\frac{n_{z}}{2}$, and ${\mathbf{Q}}_{S}$ is processed the same way as ${\mathbf{Q}}_{B}$ to obtain voxel queries. To leverage ${\mathbf{C}}_{B}$ for high-level refinement, it is downsampled with a 3D convolution of stride two and combined with new voxel encodings through element-wise addition. We get the final compact cost volume ${\mathbf{C}}\in\mathbb{R}^{\frac{n_{x}}{2}\times\frac{n_{y}}{2}\times\frac{n_{z}}{2}\times D}$ in a way similar to how ${\mathbf{C}}_{B}$ is computed.
@@ III-D Model Training
The ODTFormer is trained in a progressive manner. We first train the detection module to obtain reliable detection results. We then jointly train the detection and tracking modules together, where the feature representations of the voxels are refined to encode the spatio-temporal cues effectively.
@@ IV-B Obstacle Tracking
ODTFormer (Ours) 0.021 1.11 64.47G
We finetune ODTFormer for voxel tracking using the synthetic SceneFlow Driving dataset [42] and the real-world KITTI 2015 dataset [30], which contains over $4,000$ stereo image sequences with the resolution of $540\times 960$ and 200 stereo images with the resolution of $370\times 1224$, respectively. We apply the same data augmentation as obstacle detection and voxelize the scene flow by averaging the scene flow vectors of the point cloud within occupied voxels. We set the FPS and velocity thresholds for the matching boundary as $f=26$ FPS and $v=33.3m/s$ (or 120 $km/h$). The volumetric region of each voxel during tracking is set to be the surrounding $9\times 3\times 9$ voxels along $x$, $y$, $z$ axes.
@@
Ours w/o geometric constraints 1 6.17 9.92 77.77 64.01 23.79G 9.52M
@@
Ours w/o matching cost refinement 1 3.73 5.49 82.53 72.92 24.65G 6.22M
@@ IV-C Ablation Studies
We first replace the feature concatenation in Eq.(3) with cross-view averaging as in existing work [6, 7, 8, 9]. In this way, it discards the geometric constraints between two stereo images. As we can see at the top of Table III, it leads to significantly worse detection accuracy.
@@ V Conclusion
This paper contributed ODTFormer, a Transformer-based model to address obstacle detection and tracking problems. We detect the obstacles in the form of a voxel occupancy grid using a deformable attention mechanism and track them by matching voxels across two consecutive frames. We confirmed the effectiveness of ODTFormer by providing extensive quantitative and qualitative results on DrivingStereo and KITTI datasets. Our results showed that ODTFormer achieves state-of-the-art performance on obstacle detection tasks. On the obstacle tracking task, we showed that ODTFormer can achieve comparable accuracy to the state-of-the-art methods. The entire model is efficient and runs fast at 20fps for the KITTI resolution ($370\times 1224$) on an RTX A5000 GPU without careful postprocessing (e.g., quantization, pruning). Being able to detect and track obstacles at the same time will empower a set of downstream tasks, for instance, ensuring safe navigation when people move around a robot. We will release the code and model weights to foster future research.

### v3

@@ Abstract
Obstacle detection and tracking represent a critical component in robot autonomous navigation. In this paper, we propose ODTFormer, a Transformer-based model that addresses both obstacle detection and tracking problems. For the detection task, our approach leverages deformable attention to construct a 3D cost volume, which is decoded progressively in the form of voxel occupancy grids. We further track the obstacles by matching the voxels between consecutive frames. The entire model can be optimized in an end-to-end manner. Through extensive experiments on DrivingStereo and KITTI benchmarks, our model achieves state-of-the-art performance in the obstacle detection task. We also report comparable accuracy to state-of-the-art obstacle tracking models while requiring only a fraction of their computation cost, typically ten-fold to twenty-fold less. Our code is available on https://github.com/neu-vi/ODTFormer.
@@ I Introduction
In this paper, we present ODTFormer, a Transformer-based [23] model, to address the aforementioned obstacle detection and tracking challenges. Unlike existing stereo-based models, which construct a pixel-wise cost volume for disparity matching [24, 25, 26, 27, 14], we propose to use deformable cross-attention [28] from 3D voxel queries to 2D stereo image features to compute the matching cost. Compared with the pixel-wise cost volume used in [14], we directly construct it in the 3D space, conforming better to the scene geometry. More importantly, our approach disentangles dataset specifics from the model design and thus shows better generalization than [14]. The cost volume is then processed by a U-Net decoder to produce occupancy voxel grids progressively for efficiency purposes.
@@
To validate the effectiveness of our approach, we conduct comprehensive evaluations on DrivingStereo [29] and KITTI [30] datasets. In the obstacle detection task, we demonstrate a significant performance improvement compared to prior works [24, 25, 26, 27, 14]. Additionally, our approach exhibits greater generalization across various camera parameters and resolutions than [14]. In the obstacle tracking task, we compare scene flow-based approaches and show that our approach can achieve comparable results to state-of-the-art methods while needing much lighter computation cost. Detailed ablation studies are conducted to validate different design choices. In summary, our contributions are:
@@ II Related Works
Figure 2: Illustration of the overall architecture design. Left: For obstacle detection, we first extract multi-scale 2D feature maps [31, 32] for each of the stereo images. We then encode the voxels in the ROI to cross-attend to the image features to compute the matching cost through our novel cost volume construction method. Such a cost volume is directly constructed in the 3D space, which conforms better to the scene geometry, disentangles dataset specifics from model design, and thus generalizes well. It is then progressively decoded into occupancy voxel grids. Right: For obstacle tracking, we cast it as a matching problem by finding the correspondences of voxels across two consecutive frames, where we incorporate physical constraints to improve both the accuracy and efficiency. Both the detection and tracking modules can be jointly optimized in an end-to-end manner and run efficiently.
@@ II-B Obstacle Perception
Obstacle tracking aims at estimating the motion of obstacles across time. While most of the previous works are based on LiDAR point cloud [36, 37] or camera-LiDAR fusion [38], visual end-to-end methods, especially stereo-based and occupancy-based ones, remain under-explored. Existing visual approaches rely on object detectors or Kalman filters [10, 12, 11], limiting generalizability and requiring meticulous parameter tuning. We introduce voxel flow, which is the 3D grid-to-grid voxel-level motion of occupied voxels while achieving near real-time obstacle tracking.
In parallel to our work, there exist seminal occupancy detection approaches tailored for autonomous driving [8, 7]. These approaches commonly operate under the assumption of having access to images from the surrounding views. In contrast, our approach utilizes only a pair of stereo images and emphasizes computational efficiency.
@@ III-A2 Cost volume construction via cross-attention
A crucial difference of our proposed approach from [14] is our novel cost volume construction module design, where we employ cross-attention from 3D voxel queries to 2D stereo image features. Compared with the pixel-wise cost volume used in [14], we directly construct it in the 3D space, conforming better to the scene geometry. More importantly, our approach disentangles dataset specifics from the model design and thus shows better generalization than [14]. We will introduce this module in detail in Sec. III-B.
@@ III-B Cost Volume Construction via Cross-Attention
Specifically, for the voxel centered at $p$, we first sample points on multi-scale feature maps $({\mathbf{F}}_{i}^{L},{\mathbf{F}}_{i}^{R})$ (similar to the pixel-aligned features ${\mathbf{f}}$ but on multi-scale features instead of just the last level) around the projected points of the voxel centroid $p$ with the offset $\delta$, respectively. $\delta$ is generated by a learnable 3D offset sampler $g$ [28] based on the voxel’s encoding ${\mathbf{q}}$, $\delta=g({\mathbf{q}})$, where $g$ is an MLP plus a sigmoid output layer (scaled by the 3D voxel size of the coarsest output, see experiment settings in Sec. IV-A). We then directly concatenate the sampled features along the channel dimension and employ a scale-specific MLP to compute the cross-view matching cost ${\mathbf{c}}_{i}$:
${\mathbf{c}}_{i}(p,{\mathbf{q}},{\mathbf{F}}_{i})=\texttt{MLP}_{i}($ ${\mathbf{F}}_{i}^{L}(\mathcal{P}(p+\delta,\theta^{L}))\oplus$ (3)
${\mathbf{F}}_{i}^{R}(\mathcal{P}(p+\delta,\theta^{R}))),$
where $\oplus$ denotes the feature concatenation. $\texttt{MLP}_{i}$ has two linear layers with batch normalization and ReLU activation in between. Since we project the same 3D sample point onto multi-scale feature maps, the resultant multi-scale matching costs for this 3D point can be aggregated by concatenating along the channel dimension and employing an MLP to compute the voxel matching cost ${\mathbf{c}}$:
${\mathbf{c}}(p,{\mathbf{q}},{\mathbf{F}}_{\{1,2,3,4\}})=\texttt{MLP}($ ${\mathbf{c}}_{1}(p,{\mathbf{q}},{\mathbf{F}}_{1})\oplus{\mathbf{c}}_{2}(p,{\mathbf{q}},{\mathbf{F}}_{2})\oplus$ (4)
${\mathbf{c}}_{3}(p,{\mathbf{q}},{\mathbf{F}}_{3})\oplus{\mathbf{c}}_{4}(p,{\mathbf{q}},{\mathbf{F}}_{4})),$
where MLP has the same layer configuration as $\texttt{MLP}_{i}$. Instead of using a single offset, multiple ones are usually generated to enhance the computing of the matching cost, and different matching costs for different samples are aggregated as follows.
${\mathbf{C}}_{B}(p,{\mathbf{q}},{\mathbf{F}})=\sum_{s=1}^{N_{s}}\mathbf{A}_{s}\mathbf{W}_{s}{\mathbf{c}}(p,\delta_{s},{\mathbf{F}}_{\{1,2,3,4\}}),$ (5)
where $N_{s}$ is the total number of sampling points. $\delta_{s}$ denotes the $s$-th learned offset. $\mathbf{A}_{s}\in[0,1]$ is the learnable weights for the cost generation and $\mathbf{W}_{s}\in\mathbb{R}^{D\times D}$ are learnable parameters. We can get a cost volume ${\mathbf{C}}_{B}\in\mathbb{R}^{n_{x}\times n_{y}\times n_{z}\times D}$ by computing the matching cost for all voxels in ${\mathbf{Q}}_{B}$.
Deformable cross-attention has been used in previous works [6, 7, 8, 9] to aggregate multi-view and multi-scale features. But instead of simply averaging them as in existing work, we leverage the geometric constraints between two stereo images and use the deformable attention to construct a cost volume in 3D, leading to better accuracy, as evidenced by our ablation studies.
@@ III-D Model Training
The ODTFormer is trained progressively. We first train the detection module to obtain reliable detection results. We then jointly train the detection and tracking modules together, where the feature representations of the voxels are refined to encode the spatio-temporal cues effectively.
@@ IV-B Obstacle Tracking
ODTFormer (Ours) 0.021 1.08 64.47G
We finetune ODTFormer for voxel tracking using the synthetic SceneFlow Driving dataset [42] and the real-world KITTI 2015 dataset [30], which contains over $4,000$ stereo image sequences with the resolution of $540\times 960$ and 200 stereo images with the resolution of $370\times 1224$, respectively. We resize the images and camera parameters to have the same resolution $400\times 880$ as in DrivingStereo, then apply the same data augmentation as obstacle detection and voxelize the scene flow by averaging the scene flow vectors of the point cloud within occupied voxels. We set the FPS and velocity thresholds for the matching boundary as $f=26$ FPS and $v=33.3m/s$ (or 120 $km/h$). The volumetric region of each voxel during tracking is set to be the surrounding $9\times 3\times 9$ voxels along $x$, $y$, $z$ axes.
@@
Ours w/o geometric constraints 1 6.17 9.92 77.77 64.01 23.79G 7.52M
@@
Ours w/o 3D volumetric sampling 1 3.73 5.49 82.53 72.92 24.65G 6.22M
@@ IV-C Ablation Studies
We first replace the feature concatenations in Eq.(3) and (4) with cross-view averaging as in existing work [6, 7, 8, 9]. In this way, it discards the geometric constraints between two stereo images. As we can see at the top of Table III, it leads to significantly worse detection accuracy.
@@ V Conclusion
This paper contributed ODTFormer, a Transformer-based model to address obstacle detection and tracking problems. We detect the obstacles in the form of a voxel occupancy grid using a deformable attention mechanism and track them by matching voxels across two consecutive frames. We confirmed the effectiveness of ODTFormer by providing extensive quantitative and qualitative results on DrivingStereo and KITTI datasets. Our results showed that ODTFormer achieves state-of-the-art performance on obstacle detection tasks. For the obstacle tracking task, we showed that ODTFormer can achieve accuracy comparable to state-of-the-art methods. The entire model is efficient and runs fast at 20fps for the KITTI resolution ($370\times 1224$) on an RTX A5000 GPU without careful postprocessing (e.g., quantization, pruning). Being able to detect and track obstacles at the same time will empower a set of downstream tasks, for instance, ensuring safe navigation when people move around a robot.